Five years from now, when you’ve graduated and started your career, your salary is $65,000 per year.  During the following year, the inflation rate is 6%.  How much would your salary have to rise during that year to keep the purchasing power (or real value) of your salary from falling?  Explain, and show your work.
Answer #1

Answer :-

Here it is given that inflation rate is 6%

So the product that is of $100 at the time when the salary is $65000 then in the following the salary needs to increased at the same inflation rate by which we keep the equal purchasing power of the salary

Let us suppose a toy is of $100 when the salary was $65000 means now a person can purchase 650 toys by his salary

Then due to inflation rate the product in the following month is of $106 then to get 650 toys now the salary must be 106*650 = $68900

Means the new salary need to be = 68900/65000*100 = 106% of previous salary

Therefore salary must rise 106% - 100% = 6%

That measn equal to inflation rate